pons 5. For 1000 men age 18-24 in the HANES sample Average height (X) 75 inches SD -3 inches Average weight()-160 pounds SD, 20 pounds Correlation r .6 One man in the sample was 65 inches tall what is his expected weight? (10 points) 6. The observed ages of a group of interns at a local hospital were: 33, 37, 32, 27, 28, 36, 38). Find the median of this list. a. b. Find the mode of this list. c. Find the average and the standard deviation of this list of age. d. How many of the ages were within SD of the average age? e. How many of the ages were within 1.75 SD of the average age? f. To convert the observed value of 28 to standard units. (20 points)
Answer #1

5 ) Mean of X= 75

Mean of Y= 160

Sx= 3

Sy= 20

r= 0.6

Y'= b0+ b1*X

b_1= r*(\frac{S_y}{S_x})=0.6*(\frac{20}{3})=4

b_0= \bar{Y}-b_1*\bar{X}=160-(4*75)=-140

Y'=-140+4*X

X= 65

Y'=-140+4*65=-140+260=120